### v1

@@ 1 Introduction
On the theoretical side, various models have been employed to explore the GFFs of pNGBs [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] and nucleons [41, 42, 43, 44, 45, 46, 25, 47, 26, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. However, these model-dependent results are typically subject to uncertainties that are difficult to control. Lattice QCD (LQCD) has been used to compute the quark contribution to the GFFs of the pion and proton [60, 61, 62, 63, 64, 65], the gluonic contribution to the GFFs for various hadrons [66, 67], and the gluonic trace anomaly FFs of the pion and nucleon [68]. A notable development is the recent LQCD calculations of the pion and nucleon GFFs [69, 70] with a pion mass of $m_{\pi}=170$ MeV, which is close to the physical pion mass. Both the light quark and gluonic contributions are considered therein [69, 70], leading to significant improvement compared to earlier LQCD calculations. Recently, we have computed the pion and nucleon GFFs using the technique of dispersion relation (DR) [71], and obtained model-independant results in the sense that the method is data driven based on axiomatic principles such as unitarity, analyticity and crossing symmetry (see, e.g. Ref. [72], for a recent comprehensive review) and the uncertainties are under control.
While mass and spin are related to the Casimir operators of the Poincaré group, the nucleon $D$-term is associated with the stress component of the EMT, which reveals the mechanical properties of the nucleon (for a different point of view, see Ref. [73]). However, it remains poorly understood and is therefore regarded as “the last global unknown property” of the nucleon [2, 3]. Unlike the case of pions, the value of the nucleon $D$-term is not directly determined by chiral symmetry, but is indirectly influenced by the complex internal dynamics of QCD [74]. Based on classical mechanical stability arguments, it is conjectured that the $D$-term should be negative [21, 75]. This conjecture is supported by various chiral soliton model calculations [76, 77, 78, 79, 80, 81]: $-4\lesssim D\lesssim-1$ for the nucleon, and also by the result from dispersive analysis in Ref. [71], $D=-3.38_{-0.35}^{+0.34}$.
@@
The outline of this paper is as follows. In Sec. 2, we briefly introduce the EMT and the definition of GFFs. In Sec. 3, we review our framework in Ref. [71] for analyzing the GFFs of pNGBs by considering the $\pi\pi$ and $K\bar{K}$ intermediate states and the corresponding unitarity relations, matched to the NLO ChPT at low momentum transfer; then we present results at a few unphysical pion masses up to 391 MeV. The framework for a comprehensive description of the nucleon GFFs is reviewed in Sec. 4.1. In Sec. 4.2, we show our results of various internal spatial distributions and the corresponding radii of the nucleons. The matching of the DR representation to the ChPT result for the nucleon will be discussed in Sec. 4.3. The results of our work are summarized in Sec. 5. Details on the use of the $z$-expansion method for the extrapolation of the GFFs to asymptotically large momentum transfer are provided in Appendix A.
@@ 2.1 Generic symmetric EMT in field theory
Noether’s first theorem relates continuous global symmetry transformations to conserved currents. Consequently, the invariance under infinitesimal spacetime transformations $x^{\mu}\rightarrow x^{\prime\mu}=x^{\mu}+\varepsilon^{\mu}$ of a system, described by a Lagrangian $\cal{L}$, leads to the canonical EMT,
@@
where $\phi$ is a real scalar field. Poincaré invariance ensures that the canonical EMT is conserved (i.e., $\partial_{\mu}T_{\text{can}}^{\mu\nu}=0$). It is symmetric (i.e., $T_{\text{can}}^{\mu\nu}=T_{\text{can}}^{\nu\mu}$) for spin-0 scalar fields, but asymmetric for higher-spin fields. The Belinfante improvement addresses this issue by adding a divergence term $\partial_{\lambda}X^{\lambda\mu\nu}$ satisfying $X^{\lambda\mu\nu}=-X^{\mu\lambda\nu}$, so that current conservation is preserved while the Lorentz indices become symmetric. Furthermore, Noether’s second theorem implies that, by carefully considering local spacetime translations and field transformation properties, a symmetric and gauge-invariant (in gauge theories) EMT can be obtained (see Ref. [82] and references therein).
@@
where $g$ denotes the determinant of the metric tensor $g^{\mu\nu}$, and $\eta^{\mu\nu}$ is the metric tensor in flat spacetime. A pedagogical description of this method can be found in Appendix E of Ref. [83]. In this work, we utilize the symmetric EMT (for a discussion of the asymmetric EMT, see Ref. [84]).
@@ 2.3 Definition of gravitational form factors
$\left\{\begin{array}[]{c}X^{s}=\frac{1}{2}\left(X^{p}+X^{n}\right)\\ X^{v}=\frac{1}{2}\left(X^{p}-X^{n}\right)\end{array}\right.\ ,\quad\left\{\begin{array}[]{c}X^{p}=X^{s}+X^{v}\\ X^{n}=X^{s}-X^{v}\end{array}\right.\ .$ (17)
@@ 2.4 Properties of gravitational form factors
    where $-hR(x)\phi^{2}/2$ is a non-minimal coupling term, and $R$ is the Ricci curvature scalar. In fact, this term does not violate any symmetry of the free scalar field and is also a renormalizable dimension-4 operator. Therefore, the GFF $D(t)$ and the $D$-term are $D(t)=D=-1+4h$, whose value is unknown as $h$ is free. If we make a trivial choice $h=0$ (also called the minimally coupled case), it results in $D=-1$ [90, 30]. However, the choice of $h$ is not unique and actually relies on additional symmetry and renormalization requirements of the theory.
  • 2)2)2 )
@@
[…]
  • 3)3)3 )
@@
    In the $D$-term, the scale dependence in the low-energy constants (LECs) $L_{11}^{r}(\mu)$ and $L_{13}^{r}(\mu)$ cancels out that in the chiral logarithms $I_{i}$. In Ref. [91], the LECs were estimated using DR techniques and the scalar meson dominance model as
@@ 3.1 Unitarity and spectral function
$\operatorname{Disc}\left\langle\pi^{a}(p^{\prime})\pi^{b}\left(p\right)\left|\hat{T}^{\mu\nu}(0)\right|0\right\rangle$
=\ = $\frac{\delta^{ab}}{2}\left[\operatorname{Disc}A^{\pi}(t)\Delta^{\mu}\Delta^{\nu}+\operatorname{Disc}D^{\pi}(t)\left(P^{\mu}P^{\nu}-tg^{\mu\nu}\right)\right]$
=\ = $2i\frac{2p_{\pi}}{\sqrt{t}}\frac{\delta^{ab}}{2}\bigg{[}\left(A^{\pi}(t)\right)^{*}\bigg{(}\frac{4}{3t}p_{\pi}^{2}(t^{0}_{0}(t)-t^{0}_{2}(t))\left(P^{\mu}P^{\nu}-tg^{\mu\nu}\right)+t^{0}_{2}(t)\Delta^{\mu}\Delta^{\nu}\bigg{)}$
@@ 3.2 Coupled-channel Muskhelishvili-Omnès formalism
In Ref. [71], the GFF $A^{\pi}$ in the single-channel case, i.e., (31), has already been thoroughly analyzed using the Omnès formalism. Building upon the methodology developed for the trace GFF analysis in [71], we extend the analysis to the $D$-wave coupled-channel case.
First, we need to provide the $D$-wave Omnès matrix in the coupled-channel scenario, i.e., (45). The $D$-wave Omnès matrix corresponds to the so-called Muskhelishvili-Omnès (MO) solution [104, 105] to the following homogeneous integral equation:
@@
[…]
@@
[…]
@@
where $m_{f_{2}}$ is the mass of the lowest-lying tensor meson $f_{2}(1270)$, which leads to $L_{12}^{r}=-F_{\pi}^{2}/(2m_{f_{2}}^{2})$. It should be noted that slightly different values have been reported for the $f_{2}(1270)$ mass, namely $(1259\pm 4\pm 4)~{}$MeV [110], $(1263\pm 12)~{}$MeV [111], $(1275\pm 6)~{}$MeV [112], and the averaged value in the Review of Particle Physics, $m_{f_{2}}=(1275.4\pm 0.8)~{}$MeV [100]. Here, we set $m_{f_{2}}=(1275\pm 20)~{}$MeV to cover all these values for error estimation.
Figure 3: The total GFFs $A^{\pi/K}$ from single-channel Omnès solution [71] and coupled-channel MO solution. The red lines show the NLO ChPT prediction in the small-$|t|$ region. We also show the LQCD results at $m_{\pi}=170~{}$MeV in Ref. [69] for comparison.
@@
[…]
@@
Figure 4: The total GFFs $A^{\pi/K}$ of the pion and kaon from Ref. [71]. The red lines show the NLO ChPT prediction in small-$|t|$ region. The LQCD results for $\Theta^{\pi}$ at $m_{\pi}=170~{}$MeV are obtained from a linear combination of $A^{\pi}$ and $D^{\pi}$ in Ref. [69], with errors added in quadrature.
@@
Figure 5: The total GFFs $D^{\pi/K}$ of the pion and kaon. The red lines show the NLO ChPT prediction in the small-$|t|$ region. We also show the LQCD data at $m_{\pi}=170~{}$MeV in Ref. [69] for comparison.
@@ 3.3 Pion mass dependence of the pion GFFs
The pion mass dependence has been studied in the literature for the pion vector FF [115], which is connected to the corresponding $P$-wave $\pi\pi$ scattering amplitudes using the Omnès representation. Here, we aim to present the pion mass dependence of the pion GFFs. For the trace GFF $\Theta^{\pi}$, we rely on the prediction of $\dot{\Theta}^{\pi}(0)$ from $\mathcal{O}(p^{4})$ ChPT and the $S$-wave phase shifts at unphysical pion masses from crossing-symmetric Roy-equation analyses [116, 117]. However, for the tensor GFF $A^{\pi}$, given the lack of precise $D$-wave phase shifts at unphysical pion masses, we adopt the TMD model in Eq. (62) instead.
@@
For $m_{\pi}=239$ and 283283283~{}283MeV, we choose the matching point between the input phase shift and the extrapolation at $E_{0}\simeq 0.9$ GeV, which is below the $f_{0}(980)$ and the $K\bar{K}$ threshold. Similarly to Eq. (55), the extrapolation to the asymptotic value $\pi$ at $t\rightarrow+\infty$ is done via
@@
The TMD model adequately describes the GFF $A^{\pi}$ [122, 37] (cf. Eq. (62)), given that the $I=0$ $D$-wave $\pi\pi$ scattering is primarily governed by the $f_{2}(1270)$ resonance. Here, we use the latest prediction for the pion mass dependence of $m_{f_{2}}$ from the resonance chiral theory approach [123]. Given the small error associated with this prediction, we add a fixed uncertainty of ±$\pm 20$ MeV to $m_{f_{2}}$, as in the physical case. The calculated results are presented in Fig. 6. The variation of $A^{\pi}$ is quite mild as $m_{\pi}$ increases. By applying Eq. (37), we can also determine the behavior of $D^{\pi}$ as $m_{\pi}$ varies. Figure 6 shows that $D^{\pi}$ undergoes a significant change when $m_{\pi}$ increases from 283 to 391 MeV, which is inherited from that of $\Theta^{\pi}(t)$. Thus, we can conclude that the mechanical properties of the pion exhibit minor changes for small pion masses. However, significant changes emerge when $m_{\pi}$ is large enough for the $\sigma$ meson to become a bound state.

### v2

@@ 1 Introduction
On the theoretical side, various models have been employed to explore the GFFs of pNGBs [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] and nucleons [42, 43, 44, 45, 46, 47, 25, 48, 26, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. However, these model-dependent results are typically subject to uncertainties that are difficult to control. Lattice QCD (LQCD) has been used to compute the quark contribution to the GFFs of the pion and proton [61, 62, 63, 64, 65, 66, 67], the gluonic contribution to the GFFs for various hadrons [68, 69], and the gluonic trace anomaly FFs of the pion and nucleon [70]. A notable development is the recent LQCD calculations of the pion and nucleon GFFs [71, 72] with a pion mass of $m_{\pi}=170$ MeV, which is close to the physical pion mass. Both the light quark and gluonic contributions are considered therein [71, 72], leading to significant improvement compared to earlier LQCD calculations. Recently, we have computed the pion and nucleon GFFs using the technique of dispersion relation (DR) [73], and obtained model-independant results in the sense that the method is data driven based on axiomatic principles such as unitarity, analyticity and crossing symmetry (see, e.g. Ref. [74], for a recent comprehensive review) and the uncertainties are under control.
While mass and spin are related to the Casimir operators of the Poincaré group, the nucleon $D$-term is associated with the stress component of the EMT, which reveals the mechanical properties of the nucleon (for a different point of view, see Ref. [75]). However, it remains poorly understood and is therefore regarded as “the last global unknown property” of the nucleon [2, 3]. Unlike the case of pions, the value of the nucleon $D$-term is not directly determined by chiral symmetry, but is indirectly influenced by the complex internal dynamics of QCD [76]. Based on classical mechanical stability arguments, it is conjectured that the $D$-term should be negative [21, 77]. This conjecture is supported by various chiral soliton model calculations [78, 79, 80, 81, 82, 83]: $4\lesssim D\lesssim-1$ for the nucleon, and also by the result from dispersive analysis in Ref. [73], $D=-3.38_{-0.35}^{+0.34}$.
@@
The outline of this paper is as follows. In Sec. 2, we briefly introduce the EMT and the definition of GFFs. In Sec. 3, we review our framework in Ref. [73] for analyzing the GFFs of pNGBs by considering the $\pi\pi$ and $K\bar{K}$ intermediate states and the corresponding unitarity relations, matched to the NLO ChPT at low momentum transfer; then we present results at a few unphysical pion masses up to 391 MeV. The framework for a comprehensive description of the nucleon GFFs is reviewed in Sec. 4.1. In Sec. 4.2, we show our results of various internal spatial distributions and the corresponding radii of the nucleons. The matching of the DR representation to the ChPT result for the nucleon will be discussed in Sec. 4.3. The results of our work are summarized in Sec. 5. Details on the use of the $z$-expansion method for the extrapolation of the GFFs to asymptotically large momentum transfer are provided in Appendix B.
@@ 2.1 Generic symmetric EMT in field theory
[…]
@@
where $\phi$ is a real scalar field. Poincaré invariance ensures that the canonical EMT is conserved (i.e., $\partial_{\mu}T_{\text{can}}^{\mu\nu}=0$). It is symmetric (i.e., $T_{\text{can}}^{\mu\nu}=T_{\text{can}}^{\nu\mu}$) for spin-0 scalar fields, but asymmetric for higher-spin fields. The Belinfante improvement addresses this issue by adding a divergence term $\partial_{\lambda}X^{\lambda\mu\nu}$ satisfying $X^{\lambda\mu\nu}=-X^{\mu\lambda\nu}$, so that current conservation is preserved while the Lorentz indices become symmetric.
@@
where $g$ denotes the determinant of the metric tensor $g^{\mu\nu}$, and $\eta^{\mu\nu}$ is the metric tensor in flat spacetime. A pedagogical description of this method can be found in Appendix E of Ref. [84]. In this work, we utilize the symmetric EMT (for a discussion of the asymmetric EMT, see Refs. [85, 86] and references therein).
@@ 2.3 Definition of gravitational form factors
$\left\{\begin{array}[]{c}X^{s}=\frac{1}{2}\left(X^{p}+X^{n}\right)\\ X^{v}=\frac{1}{2}\left(X^{p}-X^{n}\right)\end{array}\right.\ ,\quad\left\{\begin{array}[]{c}X^{p}=X^{s}+X^{v}\\ X^{n}=X^{s}-X^{v}\end{array}\right.\$. (17)
@@ 2.4 Properties of gravitational form factors
    where $hR(x)\phi^{2}/2$ is a non-minimal coupling term, and $R$ is the Ricci curvature scalar. In fact, this term does not violate any symmetry of the free scalar field and is also a renormalizable dimension-4 operator. Therefore, the GFF $D(t)$ and the $D$-term are D(t)=D=1+4hD(t)=D=-1+4h, whose value is unknown as $h$ is free. If we make a trivial choice $h=0$ (also called the minimally coupled case), it results in D=1D=-1 [92, 31]. However, the choice of $h$ is not unique and actually relies on additional symmetry and renormalization requirements of the theory.
  • $2)$
@@
    Due to the presence of the four-particle $\phi^{4}$ interaction, one-loop renormalizability requires h=1/6h=1/6333This is also called the conformally coupled case [89], which means that if $m=0$ the action and hence the field equations are invariant under conformal transformations. and then $D=-1+4\times 1/6=-1/3$ [89]; see also Refs. [31, 99]. To be more specific, the choice $h=1/6$ removes UV divergences of $D(t)$ up to three loops in dimensional regularization in $\phi^{4}$ theory, although the value of $h$ needs to be changed beyond three loops [100].
  • $3)$
@@
    In the $D$-term, the scale dependence in the low-energy constants (LECs) $L_{11}^{r}(\mu)$ and $L_{13}^{r}(\mu)$ cancels out that in the chiral logarithms $I_{i}$. In Ref. [93], the LECs were estimated using DR techniques and the scalar meson dominance model as [^4] Various chiral models can be used to estimate the curvature-induced LECs $L_{11},L_{12}$ and $L_{13}$; see Refs. [102, 103, 104]. Recently, an extraction of the above LECs from lattice QCD calculations has been carried out in Ref. [105].
@@ 3.1 Unitarity and spectral function
[…]
$=\$$2i\frac{2p_{\pi}}{\sqrt{t}}\frac{\delta^{ab}}{2}\bigg[\left(A^{\pi}(t)\right)^{*}\bigg(\frac{4}{3t}p_{\pi}^{2}(t^{0}_{0}(t)-t^{0}_{2}(t))\left(P^{\mu}P^{\nu}-tg^{\mu\nu}\right)+t^{0}_{2}(t)\Delta^{\mu}\Delta^{\nu}\bigg)$
@@ 3.2 Coupled-channel Muskhelishvili-Omnès formalism
In Ref. [73], the GFF $A^{\pi}$ in the single-channel case, i.e., Eq. (31), has already been thoroughly analyzed using the Omnès formalism. Building upon the methodology developed for the trace GFF analysis in [73], we extend the analysis to the $D$-wave coupled-channel case.
First, we need to provide the $D$-wave Omnès matrix in the coupled-channel scenario, i.e., Eq. (45). The $D$-wave Omnès matrix corresponds to the so-called Muskhelishvili-Omnès (MO) solution [110, 111] to the following homogeneous integral equation:
@@
$\begin{aligned} &A^{\pi}(t)=\left(1+\alpha_{\pi}t\right)\left(\bm{\Omega}_{2}^{0}\right)_{11}(t)+\frac{2}{\sqrt{3}}\left(1+\alpha_{K}t\right)\left(\bm{\Omega}_{2}^{0}\right)_{12}(t)\ ,\\ &A^{K}(t)=\frac{\sqrt{3}}{2}\left(1+\alpha_{\pi}t\right)\left(\bm{\Omega}_{2}^{0}\right)_{21}(t)+\left(1+\alpha_{K}t\right)\left(\bm{\Omega}_{2}^{0}\right)_{22}(t)\ ,\end{aligned}$ (59)
@@
$\begin{aligned} &\alpha_{\pi}=\dot{A}^{\pi}(0)-\left(\dot{\bm{\Omega}}_{2}^{0}\right)_{11}(0)-\frac{2}{\sqrt{3}}\left(\dot{\bm{\Omega}}_{2}^{0}\right)_{12}(0)\ ,\\ &\alpha_{K}=\dot{A}^{K}(0)-\frac{\sqrt{3}}{2}\left(\dot{\bm{\Omega}}_{2}^{0}\right)_{21}(0)-\left(\dot{\bm{\Omega}}_{2}^{0}\right)_{22}(0)\ ,\end{aligned}$ (60)
@@
where $m_{f_{2}}$ is the mass of the lowest-lying tensor meson $f_{2}(1270)$, which leads to $L_{12}^{r}=-F_{\pi}^{2}/(2m_{f_{2}}^{2})$. It should be noted that slightly different values have been reported for the $f_{2}(1270)$ mass, namely $(1259\pm 4\pm 4)$~MeV [116], $(1263\pm 12)$~MeV [117], $(1275\pm 6)$~MeV [118], and the averaged value in the Review of Particle Physics, $m_{f_{2}}=(1275.4\pm 0.8)$~MeV [106]. Here, we set $m_{f_{2}}=(1275\pm 20)$~MeV to cover all these values for error estimation. For illustration, we show the spectral function $\operatorname{Im}A^{\pi}$ in Fig. 3 for the coupled-channel case (59), with the single-channel analogue from Ref. [73] also shown for comparison.
Figure 3: Results of the spectral function $\operatorname{Im}A^{\pi}(t)$.
Figure 4: The total GFFs $A^{\pi/K}$ from single-channel Omnès solution [73] and coupled-channel MO solution. The blue dashed lines show the NLO ChPT prediction in the small-$|t|$ region. We also show the LQCD results at $m_{\pi}=170$~MeV in Ref. [71] for comparison.
@@
$\begin{aligned} &\beta_{\pi}=\dot{\Theta}^{\pi}(0)-2m_{\pi}^{2}\left(\dot{\bm{\Omega}}_{0}^{0}\right)_{11}(0)-\frac{4m_{K}^{2}}{\sqrt{3}}\left(\dot{\bm{\Omega}}_{0}^{0}\right)_{12}(0)\ ,\\ &\beta_{K}=\dot{\Theta}^{K}(0)-\sqrt{3}m_{\pi}^{2}\left(\dot{\bm{\Omega}}_{0}^{0}\right)_{21}(0)-2m_{K}^{2}\left(\dot{\bm{\Omega}}_{0}^{0}\right)_{22}(0)\ ,\end{aligned}$ (66)
@@
The spectral function $\operatorname{Im}\Theta^{\pi}(t)$ is shown in Fig. 5. In the vicinity of the $K\bar{K}$ threshold, where the scalar $f_{0}(980)$ resonance is located, $\operatorname{Im}\Theta^{\pi}(t)$ drops sharply below zero, consistent with previous analyses [15, 38]. One sees that from 1.2 to 2.2 GeV, $\operatorname{Im}\Theta^{\pi}(t)$ does not exhibit the alternating positive and negative pattern observed in Ref. [38]. This is primarily because we employ a simple extrapolation similar to Eq. (55) for the $S$-wave phase shift above 1.3 GeV, where the $\pi\pi,K\bar{K}$ coupled-channel approximation becomes invalid. For a detailed analysis of the high-energy behavior of the spectral function of the GFF $\Theta^{\pi}$, see Refs. [38, 119].
@@
Figure 5: Results of the spectral function $\operatorname{Im}\Theta^{\pi}(t)$. The blue dashed line shows the NLO ChPT prediction in small-$|t|$ region.
Figure 6: The total GFFs $A^{\pi/K}$ of the pion and kaon from Ref. [73]. The blue dashed lines show the NLO ChPT prediction in small-$|t|$ region. The LQCD results for $\Theta^{\pi}$ at $m_{\pi}=170$~MeV are obtained from a linear combination of $A^{\pi}$ and $D^{\pi}$ in Ref. [71], with errors added in quadrature.
@@
Figure 7: The total GFFs $D^{\pi/K}$ of the pion and kaon. The blue dashed lines show the NLO ChPT prediction in the small-$|t|$ region. We also show the LQCD data at $m_{\pi}=170$~MeV in Ref. [71] for comparison.
@@ 3.3 Pion mass dependence of the pion GFFs
The pion mass dependence has been studied in the literature for the pion vector FF [122], which is connected to the corresponding $P$-wave $\pi\pi$ scattering amplitudes using the Omnès representation. Here, we aim to present the pion mass dependence of the pion GFFs. For the trace GFF $\Theta^{\pi}$, we rely on the prediction of $\dot{\Theta}^{\pi}(0)$ from $\mathcal{O}(p^{4})$ ChPT and the $S$-wave phase shifts at unphysical pion masses from crossing-symmetric Roy-equation analyses [123, 124]. The Roy-equation formalism is model-independent and does not depend on the unitarization method as used in Ref. [122] for the pion vector FF. However, for the tensor GFF $A^{\pi}$, given the lack of precise $D$-wave phase shifts at unphysical pion masses, we adopt the TMD model in Eq. (62) instead. We also notice that the pion mass effects have been recently discussed in the pion GFF case [38, 105] using ChPT with the $\overline{\rm MS}$ scheme to compare with the lattice QCD results at $m_{\pi}=170$ MeV [71].
@@
For $m_{\pi}=239$ and 283283~MeV, we choose the matching point between the input phase shift and the extrapolation at $E_{0}\simeq 0.9$ GeV, which is below the $f_{0}(980)$ and the $K\bar{K}$ threshold. Similarly to Eq. (55), the extrapolation to the asymptotic value $\pi$ at $t\rightarrow+\infty$ is done via
@@
The TMD model adequately describes the GFF $A^{\pi}$ [129, 38] (cf. Eq. (62)), given that the $I=0$ $D$-wave $\pi\pi$ scattering is primarily governed by the $f_{2}(1270)$ resonance. Here, we use the latest prediction for the pion mass dependence of $m_{f_{2}}$ from the resonance chiral theory approach [130]. Given the small error associated with this prediction, we add a fixed uncertainty of $\pm 20$ MeV to $m_{f_{2}}$, as in the physical case. The calculated results are presented in Fig. 8. The variation of $A^{\pi}$ is quite mild as $m_{\pi}$ increases. By applying Eq. (37), we can also determine the behavior of $D^{\pi}$ as $m_{\pi}$ varies. Figure 8 shows that $D^{\pi}$ undergoes a significant change when $m_{\pi}$ increases from 283 to 391 MeV, which is inherited from that of $\Theta^{\pi}(t)$. Thus, we can conclude that the mechanical properties of the pion exhibit minor changes for small pion masses. However, significant changes emerge when $m_{\pi}$ is large enough for the $\sigma$ meson to become a bound state.